OpenStudy (anonymous):

Given the function T(z) = 7 – z, find T(–4). I just want to know if my answer is correct! I got 3.

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

7-(-4)

OpenStudy (anonymous):

Negative * Negative = Positive.

OpenStudy (anonymous):

11?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you so much! would you mind looking at one more?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

If f(x) varies directly with x and f(x) = 3 when x = 12, then what is the value of f(x) when x = 40? 4?

OpenStudy (anonymous):

\[3 = 12k \implies k = 3/12 \implies k = 1/4~~~f(x) = x/4\]